• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name JBarber

    Log entry time 03:56:30 on July30,2005

    Entry number 150068

    keyword=pressure spike

    The pressure and temperature spiked (~235 psi), setting off some alarms. Beam was shut off. everything returned to normal.


    FIGURE 1